OpenStudy (anonymous):

PLEaseEee help! Select the equations that are parallel and perpendicular to y = x + 5 and that pass through the point (-2, -1). parallel: y = -x - 1 perpendicular: y = x + 2 parallel: y = x - 1 perpendicular: y = -x + 1 parallel: y = x + 1 perpendicular: y = -x - 3 parallel: y = 2x - 2 perpendicular: y = -2x - 1

OpenStudy (jdoe0001):

what's the slope of say " y = x + 5 " anyway?

OpenStudy (anonymous):

Um, im not exactly sure..

OpenStudy (anonymous):

-x+y-5=0?

OpenStudy (jdoe0001):

\(\large \bf y = x + 5\implies y = {\color{blue}{ 1}}x + 5\) see the slope now?

OpenStudy (anonymous):

im gonna say its 1 haha

OpenStudy (jdoe0001):

hehe right or 1/1 = 1 rise/run meaning rise =1 run =1 so any parallel line to that one.. will also have the same slope of 1 so to find an equation of a parallel line to it, you're looking for an equation whose slope is 1 and passes through (-2, -1) \(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ -2}}\quad ,&{\color{blue}{ -1}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= 1 \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\) ------------------------------------------------------ now...a line that's perpendicular to y = x + 5, will have the NEGATIVE RECIPROCAL slope of it so y = x + 5 has a slope of 1 or 1/1 then \(\bf \cfrac{{\color{blue}{ 1}}}{1}\qquad negative\implies -\cfrac{{\color{blue}{ 1}}}{1}\qquad reciprocal\implies -\cfrac{1}{{\color{blue}{ 1}}}\implies -1\) so you're looking for the equation of a line that has a slope of -1 and passes through (-2, -1) \(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ -2}}\quad ,&{\color{blue}{ -1}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= -1 \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (anonymous):

Woah you really know you're stuff :O sadly i do not and im still confused LOL

OpenStudy (anonymous):

hold on let me try and solve for y

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ -2}}\quad ,&{\color{blue}{ -1}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= 1 \\ \quad \\ y-{\color{blue}{ (-1)}}={\color{green}{ 1}}(x-{\color{red}{ (-2)}})\implies y+1=1(x+2)\implies y+1=x+2\\ \qquad \uparrow\\ \textit{point-slope form} \\ \quad \\ y=?\)

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ -2}}\quad ,&{\color{blue}{ -1}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= -1 \\ \quad \\ y-{\color{blue}{ (-1)}}={\color{green}{ -1}}(x-{\color{red}{ (-2)}})\implies y+1=-1(x+2)\implies y+1=-x-2\\ \qquad \uparrow\\ \textit{point-slope form} \\ \quad \\ y=?\impliedby perpendicular\)

OpenStudy (anonymous):

oh my um um i feel so silly that im still confused

OpenStudy (jdoe0001):

well... once you plug in the values... .is just a matter of linear simplification, which the exercise assumes you know already

OpenStudy (anonymous):

for the first would it be y=x+1

OpenStudy (jdoe0001):

yeap thus " y=x+1 " is parallel to y = x + 5

OpenStudy (anonymous):

Oh yay! thank you so much! you think it would be bothersome to help me with one more?

OpenStudy (jdoe0001):

you can always post anew... thus if I dunno, someone else may, and we can revise each other